Problem 20.64

A solenoid contains 780coils of very thin wire evenly wrapped over a length of 14.0cm . Each coil is 2.50cm in diameter.

Part A

If this solenoid carries a current of 6.70A, what is the magnetic field at its center?

B= T

SubmitMy Answers
Answer #1

B = ?o n I,
where,
B = magnetic field inside the solenoid,
n = N / length = number of turns per unit length of the solenoid
I = current

The permeability of free space is 1.25664 X 10^(-6) N/A^2

=> B =0.0469T
Answer #2

inductance L = uoN^2A/l

L = 4pi*10^-7 * 780*780 * 3.14* 0.0125^2 /0.14

L = 2.677*10^-3 H

B at the centre of solonoid is B = uoNi/L

so B = 4pi*10^-7 * 780 * 6.7/0.14

B = 0.04688 T
Answer #3

B = ?o n I,
where,

n = N / L

permeability =1.25664 X 10^(-6) N/A^2

=> B =0.0464T